OpenStudy (anonymous):

Assume A dividesB, and A divides C. Determine if the following statement is true or false and prove your conjecture: A divides B +C

geerky42 (geerky42):

Since \(A\) divides \(B\) and \(A\) divides \(C\), that mean there exist integer \(m\) and \(n\) such that \(An = B\) and \(Am = C\). \(B + C = An + Am = A(n+m)\) therefore \(A\) divides \(B+C \)

OpenStudy (anonymous):

Took me time to understand your response but now after reading a few times i get it. Thanks!

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Welcome.

OpenStudy (anonymous):

I have a similar problem, so will the answer be similar? Assume A divides B, and A divides C. Determine if the following statement is true or false and prove your conjecture: A divides BC BC=An*Am=A(n+m) therefore A divides B+C?

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

Can I use numbers instead of A, B and C to prove?

geerky42 (geerky42):

Numbers won't really prove anything. All they do are showing that it is true for certain numbers A, B, and C. You need to prove that it is true for all possible integers. Yeah, process is similar. You just go like "\(BC=Am\cdot An=A(mn)\) therefore \(A\) divides \(BC\)"

OpenStudy (anonymous):

Thanks! @geerky42